# Improving Code Reviewer Recommendation: Accuracy, Latency, Workload, and Bystanders

### v1

@@ 10. Concluding Remarks
    RevRecV2: In the A/B test, we saw a top 3 accuracy of 73% a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 has been rolled out to 100% of engineers at Meta.
@@
    BystanderRecRnd: The bystander effect occurs in code review when only a team is assigned to a review a diff instead of an individual. At Meta we randomly assigned one of the top3 experts to review the diff and found a drop of -11.6% in time-in-review with no regressions in our guardrail metrics. We rolled BystanderRecRnd out to 100% of engineers.
We hope that our works will inspire other engineers, researchers, and companies to look beyond experience based recommendation to consider author-reviewer familiarity, reviewer workload, and the bystander effect of only recommending teams of reviewers.

### v2

@@ 10. Conclusion
    RevRecV2: In the A/B test, we saw a Top3 accuracy of 73%, a 14 point improvement over the previous model. We also saw a 14 times reduction in latency at p90. The recommendations were selected by authors 27% more often with the new model. We did not see any statistically significant changes in the review cycle time or the reviewer viewing time. RevRecV2 has been rolled out to 100% of engineers at Meta.
@@
    BystanderRecRnd: The bystander effect occurs in code review when only a team is assigned to review a diff instead of an individual. At Meta we randomly assigned one of the top three experts to review the diff and found a drop of -11.6% in TimeInReview with no regressions in our guardrail metrics. We rolled BystanderRecRnd out to 100% of engineers.
Our contributions to the general field of software engineering are as follows:
  1. (1)
    We updated Al-Zubaidi et al.’s (Al-Zubaidi et al., 2020) 2020 literature review. We developed a recommender based on measures from each of their categories. Interestingly, we found that reviewer relationships, i.e. how often a reviewer had reviewed an author in the past, was the strongest predictor. Expertise was less important.
  2. (2)
    We noted a large discrepancy between backtesting (the typical way code review tools are evaluated in research literature) and A/B test. While, obviously, not in all instances A/B test can be performed, future studies may discover ways to do backtesting in a way that is more reflective of what would be obtained via A/B test. In many companies code review is implemented as a service, so A/B test is feasible given that privacy and other ethical considerations are adhered to.
  3. (3)
    We described a set of goal and guardrail metrics that can be used in code reviewer recommendation. We feel that TimeSpent is particularly important because it allows researchers to ensure that reviewers are not rubberstamping reviews without thoroughly considering the code. The metric can be easily captured using tools that capture the time spent on a webpage.
  4. (4)
    We are unaware of any works that reported on the impact of latency on reviewer recommendation. When our slowest 10% of recommendations (the 90th percentile recommendation time) went from 4.4 seconds to 300ms, we saw a 27% increase in the number of recommendations that were clicked on. We suggest that future work should report how long it takes to generate a recommendation, if the 90th percentile is above 300ms, then it is likely that the suggestion will load after the author has already selected other reviewers.
  5. (5)
    Although prior works attempted to balance reviewer workload, these works did not capture the change in workload, instead measuring other variables like TimeInReview. We suggested a simple measure of workload and showed that some degree of workload balancing can work. We hope that Microsoft, Ericsson, and other companies will report on the impact of workload balancing.
  6. (6)
    We are unaware of any works that quantify the impact of assigning a team rather than an individual. We were impressed how a simple random assignment of one of top recommended reviewers could effectively mitigate the bystander effect. This finding has implications beyond software firms and may impact how tasks are assigned generally in organizations. We look forward to future work on task assignments.